Solve. Let f(x) = x2 - 8x + 16. Find x so that f(x) = 22.
Answer by ankor@dixie-net.com(22740)

Solve. Let f(x) = x^2 - 8x + 16. Find x so that f(x) = 22.
:
x^2 - 8x + 16 = 22
:
x^2 - 8x + 16 - 22 = 0
:
x^2 - 8x - 6 = 0
Use the quadratic formula to find x